OpenStudy (anonymous):

Find the diameter of a circle when its circumference is 10.

OpenStudy (anonymous):

Circumference = pi * d So if c=10 then 10 = 3.14 * d So the diameter would be 10 / pi :)

OpenStudy (campbell_st):

d = 10/pi